### v1

# Hyperelliptic Shimura curves and $L$-functions of central vanishing order at least 3

Abstract.
We use hyperelliptic Shimura curves to find triple product $L$-functions of Hilbert newforms with central vanishing orders proved to be at least 3.
1. Introduction
In [11, 26], Gross and Zagier proved that the $L$-function of a certain modular elliptic curve over $\mathbb{Q}$} has central vanishing order 3. This was essential in Goldfeld’s effective solution to the Gauss class number problem for imaginary quadratic fields [6]. According to folklore conjectures, such $L$-functions with higher central vanishing orders are very rare. Moreover, such vanishing can not be proved by numerically computation. Thus, Gross and Zagier’s example is remarkable. In this paper, we present some analogs of Gross and Zagier’s example in triple product $L$-functions of abelian varieties.
Theorem 1.1.
Let $F={\mathbb{Q}}\left(\cos\frac{1}{16}\pi\right)$. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. Let $A_{1},A_{2},A_{3}$ be 4-dimensional simple modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ and of conductor $\mathfrak{p}$}. Then
${\mathrm{ord}}_{s=2}L\left(s,h^{1}(A_{1})\otimes h^{1}(A_{2})\otimes h^{1}(A_{3})\right)\geq 3\cdot 4^{3}.$
Indeed, there are exactly four modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ of conductor $\mathfrak{p}$} and dimension 4, which are all simple. Each of them correspond to a Galois orbit of size 4 of Hilbert newforms over $F$. Then the fact that the triple product $L$-function has a holomorphic continuation to $s\in{\mathbb{C}}$ follows from [5, 17].
We actually prove a slightly more general version of 1.1 in terms of Hilbert newforms over $F$. In 7.2, we show that for any 3 of the 16 Hilbert newforms (in the four Galois orbits of size 4) above, their triple product $L$-function has central vanishing order at least 3.
The proof is similar to the one of Gross and Zagier in spirit. They used their celebrated formula on modular curves, which relates $L$-functions of modular forms to the heights of (Heegner) CM 0-cycles. We use an analog proved by Yuan, S. Zhang, and W. Zhang [24] for some Shimura curves, which relates triples product $L$-functions to the heights of modified diagonal cycles. Both of us show that our cycles accidently vanish for some reasons.
Besides, the specific modular curve used for Gross and Zagier’s order 3 example is $X_{0}(37)$ (see [26]) and it has an exceptional hyperelliptic involution found by Ogg [16]. So does the Shimura curve used by us, which is the work of Dembélé [4]. It might be a coincidence, since the hyperellipticity is essential for our result, but it is not used in [11, 26]. However, in 4.7, we show that hyperellipticity can be used to show vanishing of CM 0-cycles. And in 6.2, we show that hyperellipticity is also related to higher central vanishing of the $L$-functions of single Hilbert newforms.
Finally, we want to remark that there are only finitely many hyperelliptic Shimura curve, as shown in [20]. This is also a facet of the expectation that $L$-functions of central vanishing order at least 3 are very rare.
Acknowledgments
The author would like to thank Peter Sarnak for raising the question of on possible analogs of Gross–Zagier’s example in higher dimensions. The author is also grateful to Yifeng Liu and Wei Zhang for many helpful discussions.
2. Triple product of abelian motives
For a smooth projective variety $V$ over a field, let ${\mathrm{Ch}}_{i}(V)$ be the Chow group of $V$ of dimension $i$ cycles with $\mathbb{Q}$}-coefficients, modulo rational equivalence. For a morphism $f:V\to U$, let $\Gamma_{f}\in{\mathrm{Ch}}_{*}(V\times U)$ be the graph of $f$.
Let $A$ be an abelian variety over a field $F$ of dimension $g$, with dual abelian variety $A^{\vee}$. Let ${\mathfrak{pr}}\in{\mathrm{Ch}}_{g}(A\times A)$ be the (2g1)(2g-1)-th Deninger–Murre projector, that is denoted by $\pi_{2g-1}$ in [3, Theorem 3.1]. For a variety $X$ over $F$, the abelian group structure of $A$ induces an abelian group structure on ${\mathrm{Mor}}(X,A)$, the set of morphisms of $F$-varieties. Then by regarding $X\times A$ as an abelian scheme over $X$, [23, Corollary A.3.2] implies that the following map is a homomorphism of abelian groups:
${\mathrm{Mor}}(X,A)$ $\to{\mathrm{Ch}}_{\dim X}(X\times A),$
$f$ $\mapsto{{\mathfrak{pr}}}\circ[\Gamma_{f}].$
Denote the unique $\mathbb{C}$}-linear extension ${\mathrm{Mor}}(X,A)_{\mathbb{C}}\to{\mathrm{Ch}}_{\dim X}(X\times A)$ by
$f\mapsto{{\mathfrak{pr}}_{f}}.$
For an idempotent $\iota\in{\mathrm{End}}(A)_{\mathbb{C}}$, by [3, Proposition 3.3], ${\mathfrak{pr}}_{\iota}$ is a projector. Moreover, if $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$, one can easily check that its image $\iota^{\dagger}$ under a Rosati involution on ${\mathrm{End}}(A)_{\mathbb{C}}$ associated with a polarization does not depend on the choice of the polarization.
Proposition 2.1.
Let the symmetric group $S_{n}$ act on $A^{n}$ by permuting the components. Assume that $\dim\iota_{*}H^{1}(A)<n$. For $z\in{\mathrm{Ch}}_{*}(A^{n})$ invariant by $S_{n}$, we have $({\mathfrak{pr}}_{\iota}^{n})_{*}z=0$.
Proof.
The proof is the same as [21, Corollary 2.2.3]. ∎
Proposition 2.2.
Let $f:X\to A$ be a morphism from a smooth projective curve over $F$ whose associated Albanese morphism to $A$ is surjective. Let $\Delta\subset X^{2}$ be the diagonal. Assume that $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$ such that $\iota^{\dagger}=\iota$ and $\dim\iota_{*}H^{1}(A)\leq 2$. Then $({\mathfrak{pr}}_{\iota}^{2})_{*}{\mathrm{Ch}}_{*}(A^{2})={\mathbb{C}}\cdot({\mathfrak{pr}}_{\iota}^{2})_{*}(f^{2})_{*}[\Delta]$, which is nonzero.
Proof.
The proof is the same as [21, Corollary 2.2.7 (1)]. ∎
Remark 2.3.
The relation between the above two propositions and [21, Corollary 2.2.3 , Corollary 2.2.7 (1)] is not just similar. Indeed, if $A$ is simple, we have an isomoprhism between the motives $(A,{\mathfrak{pr}}_{\iota})$ and $(X,\delta_{f})$ defined in [21, Section 2.4][1][1][1]In loc. cit., we used $\phi$ instead of $f$ to denote a morphism from $X$ to $A$. . This isomoprhism can be easily proved using the decomposition of $\delta_{f}$ in (the summand on the right hand side of) [23, (2.13)] and the nilpotency of Kimura [13, Proposition 7.5].
3. Shimura curves
Let $F$ be a totally real number field and ${\mathbb{A}}_{\mathrm{f}}$ its ring of finite adeles. Let $B$ be a quaternion algebra split at one infinite place $\tau:F\hookrightarrow{\mathbb{R}}$ of $F$ and division at all other infinite places of $F$. Let $B^{+}\subset B^{\times}$ be the subgroup of elements with totally positive norms, equivalently, with positive determinants in $B_{\tau}^{\times}\simeq{\mathrm{GL}}_{2}({\mathbb{R}})$. Let $\mathbb{H}$} be the complex upper half-plane, and let $B^{+}\subset{\mathrm{GL}}_{2}({\mathbb{R}})$ act on $\mathbb{H}$} by the fractional linear transformation.
For an open compact-modulo-center subgroup $U$ of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, we have the smooth compactified Shimura curve $X_{U}$ for $B^{\times}$ of level $U$ over $F$ [25, 3.1.4][2][2][2]Technically, sometimes by “Shimura curve $X_{U}$ for $B^{\times}$”, it is required that $U$ is compact. And if $U$ contains the center of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, $X_{U}$ should be refered to a Shimura curve for $PB^{\times}$. We do not make this distinction here.. The complex uniformization of $X_{U}$ via the distinguished infinite place $\tau:F\hookrightarrow{\mathbb{R}}\subset{\mathbb{C}}$ is given by
(3.1) $X_{U,\tau,{\mathbb{C}}}\simeq B^{+}\backslash{\mathbb{H}}\times B^{\times}({\mathbb{A}}_{\mathrm{f}})/U\coprod\{\text{cusps}\}.$
Here, the cusps exist if and only if $X_{U}$ is a modular curve, i.e. $F={\mathbb{Q}}$, and $B$ is the matrix algebra.
[…]
For $x\in X_{U}({\mathbb{C}})$ that is the image of a point in ${\mathbb{H}}\times G({\mathbb{A}}_{\mathrm{f}})/U$ with a nontrivial stabilizer in $B^{+}$, we call $x$ an elliptic point and let $r_{x}$ be the order of the stabilizer. There are only finitely many elliptic points on $X_{U}$. Let
${\mathfrak{e}}_{U}=\sum_{\text{elliptic }x}\left(1-\frac{1}{r_{x}}\right)x$
Let $K_{U}$ be a canonical divisor of $X_{U}$, and ${\mathcal{K}}_{U}=K_{U}+{\mathfrak{e}}_{U}$ an orbifold canonical divisor of $X_{U}$. Let ${\mathfrak{c}}_{U}$ be the divisor of the cusps of $X_{U}$, which is 0 unless $X_{U}$ is a modular curve. Let
$L_{U}={\mathcal{K}}_{U}+{\mathfrak{c}}_{U}.$
[…]
$\lambda_{U}=[L_{U}]/\deg L_{U}^{\circ}$
where $L_{U}^{\circ}$ is the restriction of $L_{U}$ to any geometrically connected component of $X_{U}$. In [23, Remark 3.2.4], we further noticed that if $X_{U}$ is a modular curve, then ${\mathfrak{c}}_{U}$ is a multiple of $\lambda_{U}$. So
(3.2) $\lambda_{U}=[{\mathcal{K}}_{U}]/\deg{\mathcal{K}}_{U}^{\circ},$
provided that $\deg{\mathcal{K}}_{U}>0$.
For an abelian variety $A$ over $F$, let ${\mathrm{Mor}}^{0}(X_{U},A)$ be the subspace of $f\in{\mathrm{Hom}}(X_{U},A)_{\mathbb{Q}}$ such that for some (equivalently for every) representative $\sum_{i}a_{i}p_{i}$ of $\lambda_{U,\overline{F}}$, where $a_{i}\in{\mathbb{Q}}$ and $p_{i}\in X_{U}(\overline{F})$, the sum of $f(p_{i})\otimes a_{i}$ over $i$’s in the group $A(\overline{k})\otimes{{\mathbb{Q}}}$ is the identity of $A$. Let
[…]
$\pi_{A}^{U}={\mathrm{Mor}}^{0}(X_{U},A).$
Let ${\mathrm{End}}(A)_{\mathbb{Q}}$ act on $\pi_{A}^{U}$ and $\pi_{A}$ from left by post-composition, and one may regard $\pi_{A}$ as a $B^{\times}({\mathbb{A}}_{\mathrm{f}})$-representation with ${\mathrm{End}}(A)_{\mathbb{Q}}$-coefficients.
Theorem 3.1 ([25, 3.2.2, 3.2.3]).
Let $A$ be a simple abelian variety over $F$ such that $\pi_{A}\neq 0$.
(1) The $\mathbb{Q}$}-algebra ${\mathrm{End}}(A)_{\mathbb{Q}}$ is a finite field extension of $\mathbb{Q}$} and ${\mathrm{End}}(A)_{\mathbb{Q}}:{\mathbb{Q}}]=\dim A$.
(2) Consider the set of pairs $(A,\iota)$ where $A$ is a simple abelian variety over $F$ such that $\pi_{A}\neq 0$ and $\iota:{\mathrm{End}}(A)_{\mathbb{Q}}\hookrightarrow{\mathbb{C}}$ is an embedding. The map $(A,\iota)\to\pi_{A,\iota}$ is a bijection to the set of finite components of automorphic representations of $B^{\times}$, whose Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is cuspidal and holomorphic of weight 2.
Remark 3.2.
(1) By (1) of the theorem, $A$ is of ${\mathrm{GL}}_{2}$-type. A simple abelian variety over $F$ of ${\mathrm{GL}}_{2}$-type with $\pi_{A}\neq 0$ for some $B$ as above is called modular (or automorphic in [25]).
(2) Since ${\mathrm{End}}(A)_{\mathbb{Q}}$ is a field so that irreducible idempontents of ${\mathrm{End}}(A)_{\mathbb{Q}}$ are in bijection with embeddings $M_{A}\hookrightarrow{\mathbb{C}}$. Moreover, they are automatically central. Thus, here we use the same notation $\iota$ as in 2.1 and $2.$
Finally in this section, we explain the relation between the divisors defined above. Below we simplify notations and let $\lambda,{\mathfrak{c}},{\mathfrak{e}}$ denote $\lambda_{U},{\mathfrak{c}}_{U},{\mathfrak{e}}_{U}$ if $U$ is clear from the context.
For $z\in{\mathrm{Ch}}_{1}(X_{U})$, by [23, Corollary 2.5.3], we have
[…]
We need another class close to $\lambda$ to discuss (modified) diagonals. Assume $\deg(K_{U})>0$. Let
$\xi=[K_{U}]/\deg K_{U}^{\circ}.$
If $\deg{\mathfrak{e}}>0$, then by (3.2),
(3.4) $\xi=\lambda\text{ if and only if }[{\mathfrak{e}}]=(\deg{\mathfrak{e}})\lambda.$
By definition, ${\mathfrak{pr}}_{f,*}\lambda=0$. Then by (3.2), we have
(3.5) ${\mathfrak{pr}}_{f,*}\xi=0\text{ if and only if }{\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0.$
4. Diagonals and hyperellipticity
Consider the $n$-th power of the Shimura curve $X_{U}$. Let
[…]
where the transition maps in the inverse limit are pushforwrds by the natural morphisms between Shimura curves. Let $B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ act on $\overline{{\mathrm{Ch}}}(n)$ (from left) by pullback. Let $\pi_{i}$ be of the form $\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2), which is a representation of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$.
Theorem 4.1 ([23, Theorem 3.3.2]).
Let $H\subset B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ be a subgroup such that $\boxtimes_{i=1}^{n}\pi_{i}$ has no nonzero $H$-invariant linear forms. Then for $z=(z_{U})_{U}\in{\mathrm{Ch}}(n)$ invariant by $H$, we have
$\left(\mathop{\otimes}_{i=1}^{n}{\mathfrak{pr}}_{f_{i}}\right)_{*}z=0$
for all $U$ and all $f_{i}\in\pi_{i}^{U}$.
Let $\Delta_{n}=\Delta_{U,n}$ be the diagonal $X_{U}^{n}$. Let
$\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=\left({\mathfrak{pr}}_{f_{1}}\otimes{\mathfrak{pr}}_{f_{2}}\otimes...\otimes{\mathfrak{pr}}_{f_{n}}\right)_{*}[\Delta_{n}]$
in ${\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times...\times A_{n}).$
Corollary 4.2 ([23, Corollary 3.3.4]).
Assume that $\boxtimes_{i=1}^{n}\pi_{i}$ has no nonzero diagonal-$B^{\times}({\mathbb{A}}_{\mathrm{f}})$-invariant linear forms, then
$\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=0$
for all $U$ and all $f_{i}\in\pi_{i}^{U}$.
Here are some easy examples for applying this corollary.
Example 4.3.
(1) If $n=1$, let $H=B^{\times}({\mathbb{A}}_{\mathrm{f}})$ and we have $\Delta_{1}]_{f_{1}}=0$ for all $f_{1}\in\pi_{1}^{U}$.
(2) If $n=2$, $H$ is the diagonal $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, and $z_{U}=[\Delta_{2}]$, then $\Delta_{2}]_{f_{1},f_{2}}=0$ if $\pi_{1}\not\simeq\pi_{2}^{\vee}$.
(3) If the product of the central characters of $\pi_{i}$’s is not trivial, then $\Delta_{n}]_{f_{1},f_{2},...,f_{n}}=0$.
Proposition 4.4.
Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then
$\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$
for all $f_{i}\in\pi_{i}^{U}$.
We prove it in the end of this section. We need some preparations.
For a subset $I$ of $\{1,2,...,n\}$, let $p_{I}$ be the projection of $X_{U}^{n}$ to the product of the components indexed by $I$.
Lemma 4.5.
(1) Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then
$\Delta_{3}]_{f_{1},f_{2},f_{3}}=-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ I=\{a,b\},J=\{c\}\end{subarray}}p_{I}^{*}[\Delta_{2}]_{f_{a},f_{b}}\cdot p_{J}^{*}{\mathfrak{pr}}_{f_{c},*}\xi$
for all $f_{i}\in\pi_{i}^{U}$.
(2) Further assume $\pi_{i}\not\simeq\pi_{j}^{\vee}$ for any $i\neq j$ in $\{1,2,3\}$. Then
$\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$
for all $f_{i}\in\pi_{i}^{U}$.
Proof.
The modified diagonal defined by Gross and Schoen [10, 27]is the algebraic cycle
(4.1) $[\Delta_{3}]_{\xi}=[\Delta_{3}]-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ |I|=2\end{subarray}}p_{I}^{*}[\Delta_{2}]\cdot p_{J}^{*}\xi+\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ |I|=1\end{subarray}}p_{I}^{*}[\Delta_{1}]\cdot p_{J}^{*}(\xi\times\xi)$
By hyperellipticity and [10, 4.5], $\Delta_{\xi}]=0$. The proposition follows from 4.3 (1)(2). ∎
By 2.1 and 3.2 (2), we have the following lemma.
Lemma 4.6.
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}$ and $f_{1}=f_{2}=f_{3}=:f$, then $\Delta_{3}]_{f,f,f}=0$.
Using the last two lemmas, we are able prove the following.
Corollary 4.7.
Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then $\xi=\lambda$, ${\mathfrak{e}}]=(\deg{\mathfrak{e}})\lambda$, ${\mathfrak{pr}}_{f,*}\xi=0$ and ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$ for every $\pi=\pi_{A,\iota}$ and every $f\in\pi^{U}$.
Proof.
By (3.3) and (3.4), the four equations in the corollary are equivalent to each other. We prove the last two of them. By (3.5), the last two are equivalent to each other for each $f$.
If $\pi$ is not self-dual so that $\pi$ has nontrivial central character, then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ by using the formulation in [25, 1.3.2] with invariant linear forms on $\pi$. So we assume that $\pi$ is self-dual and want to prove that ${\mathfrak{pr}}_{f,*}\xi=0$ instead.
We use 4.5 (1) with with $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=\pi$ and $f_{1}=f_{2}=f_{3}=:f$. Indeed, by 2.1, the left hand side of the equation in (3) of the lemma is 0. Now we analyze the terms on the right hand side, and one of them is $p_{1,2}^{*}[\Delta_{2}]_{f,f}\cdot p_{3}^{*}{\mathfrak{pr}}_{f,*}\xi,$ which is just $\Delta_{2}]_{f,f}\times{\mathfrak{pr}}_{f,*}\xi$.
Claim: there is a cycle $z$ in ${\mathrm{Ch}}_{2\dim A-1}(A\times A)$ whose intersection number with $\Delta_{2}]_{f,f}$ is nonzero. Since the cohomology class of $\xi$ is 0, the intersection of the right hand side of the equation in 4.5 (1) with $z\times[A]$ is
$([\Delta_{2}]_{f,f}\times{\mathfrak{pr}}_{f,*}\xi)\cdot(z\times[A]).$
Then the push forward by $p_{3}$ of this Chow cycle is a multiple of ${\mathfrak{pr}}_{f,*}\xi$. But it is 0 by left hand side being 0.
Now we prove the claim. Since $\pi$ is self dual, by pulling back to $C$ (as in the summand on the right hand side of [23, (2.13)]), we see that the Betti cohomology class of $\Delta_{2}]_{f,f}$ in $H^{1}(A_{\mathbb{C}})\otimes H^{1}(A_{\mathbb{C}})$, which actually is in $\iota_{*}H^{1}(A_{\mathbb{C}})\otimes\iota_{*}H^{1}(A_{\mathbb{C}})$,) is nonzero. By the nonvanishing of the intersection pairing between $H^{1}(A_{\mathbb{C}})$ and $H^{2\dim A-1}(A_{\mathbb{C}})$, and thus their tensor squares, we can find $z$ as in the claim. ∎
Remark 4.8.
Note that elliptic points are CM points and $\mathfrak{e}$} is a sum of CM cycles. Then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ gives an explicit linear relation between the images by ${\mathfrak{pr}}_{f,*}$ of the CM zero cycles on $A$. This is trivial if the images are all 0, which is the case if $L^{\prime}(1/2,\pi)=0$ by the general Gross–Zagier formula [11, 25]. (In the corollary, our $\pi$ denotes the finite part of an automorphic representation of $B^{\times}$. In this remark, $L(s,\pi)$ denotes the $L$-function of this automorphic representation of $B^{\times}$.) So we assume $L^{\prime}(1/2,\pi)\neq 0$. For example, let $X_{U}$ be the classical modular curve $X_{0}(37)$, which is famously hyperelliptic by Ogg [16]. There is a unique such $\pi$, corresponding to the newform $f$ with LMFDB label 37.2.a.a. There are two elliptic points of order 2 (resp. 3) on $X_{U}$, and let ${\mathfrak{e}}_{2}$ (resp. ${\mathfrak{e}}_{3}$) be their sum. Then ${\mathfrak{pr}}_{f,*}{\mathfrak{e}}=0$ gives
(4.2) $\frac{1}{2}{\mathfrak{pr}}_{f,*}{\mathfrak{e}}_{2}+\frac{2}{3}{\mathfrak{pr}}_{f,*}{\mathfrak{e}}_{3}=0.$
This could also be read from the modular form of weight $3/2$ corresponding to $f$, by the Gross–Kohnen–Zagier formula [8]. We leave it to the interested reader. Besides, (4.2) and the version of the Gross–Zagier formula [1, Theorem 1.1] implies that the finite part of the quadratic twist $L$-functions of $\pi$ satisfies
$\frac{2^{2}\cdot\sqrt{4\cdot 2^{2}}\cdot L^{\infty}(1/2,\pi\otimes\chi_{4})}{3^{2}\cdot\sqrt{3}\cdot L^{\infty}(1/2,\pi\otimes\chi_{3})}=\frac{16}{9}.$
One can also check that this is true up to very high precision using $L$-values read from LMFDB [14].
Proof of $4.$
It follows from 4.5 and 4.7. ∎
5. Triple product $L$-functions of central vanishing order at least 3
We keep the notations from the last section, except that from now on, the notation $\pi$ or $\pi_{i}$ will denote an automorphic representation of $B^{\times}$, instead of its finite part. We always assume that it is holomorphic of weight $2$ in the sense that its Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is a cuspidal automorphic representation holomorphic of weight $2$.
For a place $v$ of $F$, let $\epsilon(B_{v})=1$ if is $B_{v}$ is split, and let $\epsilon(B_{v})=-1$ if is $B_{v}$ is division. So if $v=\tau$, $\epsilon(B_{v})=1$ and for any othet infinite place $v$, $\epsilon(B_{v})=-1.$ Then
$\prod_{v}\epsilon(B_{v})=1.$
Here the product is over all places of $F$. Let $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})$ be the local root number of $\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v}$. A result of Prasad [18, 19] states the following.
Theorem 5.1.
If
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\},$
then it has dimension 1 and $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=\epsilon(B_{v})$.
Corollary 5.2.
If
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\}$
for all finite places $v$ of $F$, then $\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Proof.
At every infinite place $v$, $\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-1$ (see [18, Section 9]). At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. So by 5.1, we have
$\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=\prod_{v}\epsilon(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-\prod_{v}\epsilon(B_{v})=-1.$
Let $S_{B}$ be the finite set of finite places $v$ of $F$ such that $B_{v}$ is a division quaternion algebra. Below we only consider a maximal open compact subgroup $U=\prod\limits_{v\text{ finite}}U_{v}\subset PB^{\times}({\mathbb{A}}_{\mathrm{f}})$. In other words, its $v$-component is
(5.1) $U_{v}=PB_{v}^{\times}\text{ for }v\in S_{B},$
and
(5.2) $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ or }F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right),g_{v}}\rangle\text{ for }v\not\in S_{B},$
where ${\mathfrak{m}}_{F_{v}}$ is the maximal ideal of ${\mathcal{O}}_{F_{v}}$ and
(5.3) $g_{v}\in\begin{bmatrix}0&1\\ {\mathcal{O}}_{F_{v}}\backslash{\mathfrak{m}}_{F_{v}}&0.\end{bmatrix}$
If $\pi^{U}\neq 0$, then $\pi$ has trivial central character and $\dim\pi_{v}^{U_{v}}=1$ for $v$ finite. Explicitly, for $v\in S_{B}$, $\pi_{v}$ is the trivial representation of $B_{v}^{\times}$. Its Jacquet-Langlands correspondence to ${\mathrm{GL}}_{2}(F_{v})$ is the the Steinberg representation. In this cases, the root number $\epsilon(\pi_{v})=-1$. For $v\not\in S_{B}$, if $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ for $v\not\in S_{B}$, then $\pi_{v}$ is an unramified representation of ${\mathrm{GL}}_{2}(F_{v})$. If $U_{v}=F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{v}\right),g_{v}}\rangle$, then $\pi_{v}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of $F_{v}^{\times}$. In both cases, the root number $\epsilon(\pi_{v})=1$.
Lemma 5.3.
If $\pi^{U}\neq 0$, then $\epsilon(\pi)=-1$.
Proof.
At an infinite place, $\epsilon(\pi_{v})=-1$. At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. And at every infinite place $v$, as we have seen above the lemma, $\epsilon(\pi_{v})=\epsilon(B_{v})$. So we have the first of the following equations $\epsilon(\pi)=-\prod_{v}\epsilon(B_{v})=-1.$
Lemma 5.4.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$ and $v$ finite, then
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\}.$
For $\ell\neq 0$ in it and $\phi_{i}\neq 0\in\pi_{i,v}^{U_{v}}$ (both unique up to scalar as seen above), we have
(5.4) $\ell(\phi_{1}\otimes\phi_{2}\otimes\phi_{3})\neq 0.$
Proof.
If $v\in S_{B}$ so that $\pi_{v}=1$, the lemma is trivial. Otherwise, the lemma is [9, Proposition 6.1, 6.3]. ∎
Then by 5.2, we have the following
Corollary 5.5.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then $\epsilon(\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Let $\pi_{i,{\mathrm{f}}}$ denote the finite component of $\pi_{i}$. Then $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). The main theorem of this section is the following.
Theorem 5.6.
Assume
  • $U$ satisfies (5.1) and (5.2),
  • every geometrically connected component of $X_{U}$ is hyperelliptic.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then
${\mathrm{ord}}_{s=1/2}L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\geq 3.$
Proof.
The generalized Gross–Kudla conjecture [7] is proved in [24] under the first two assumptions in the theorem. By this and (5.4), the central derivative $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ is a multiple of the height of $\Delta_{3}]_{f_{1},f_{2},f_{3}}$, where Let $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, unique up to scalar. But by 4.5, $\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$. Now the theorem follows from $5.$ ∎
6. Decomposable triple product $L$-functions
With the generalized Gross–Kudla conjecture[7, 24], we have used the (modified) diagonal cycle to study triple product $L$-functions. Now we take the opposite direction, and want to understand the behavior of the diagonal cycle for a general Shimura curve predicted by triple product $L$-functions. Keep the notations from the last section, except that now we abandon (5.1) and (5.2). We analyze two cases when the triple product $L$-function is decomposable.
First, if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual, then
$L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=L(s,{\mathrm{sym}}^{3}\pi)L(s,\pi)^{2}.$
So if the root number $\epsilon(1/2,\pi)=-1$, then $L(s,\pi^{\otimes 3})$ has central vanishing order at least 2 at $s=1/2$. Thus by 5.3 and 5.5, 5.6 is trivial if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}.$
For the diagonal cycle, motivated by the generalized Gross–Kudla conjecture, we propose the following conjecture.
Conjecture 6.1.
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual and $\epsilon(1/2,\pi)=-1$, then $\Delta_{3}]_{f_{1},f_{2},f_{3}}=0$ for all $f_{i}\in\pi_{i,{\mathrm{f}}}^{U}$.
4.6 gives a special case of 6.1.
Second, if $\pi_{1}\simeq\pi_{2}^{\vee}=:\pi$, then
(6.1) $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=L(s,{\mathrm{ad}}\pi\otimes\pi_{3})L(s,\pi_{3}).$
Assume $\pi_{3}$ is self-dual with root number $\epsilon(1/2,\pi_{3})=-1$. Then $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has central vanishing order at least 3 at $s=1/2$ if and only if $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has central vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has central vanishing order at least 3 at $s=1/2.$ Then by 5.3, we have the following.
Proposition 6.2.
In the situation of 5.6, if $\pi_{1}\simeq\pi_{2}=:\pi$ (which are self-dual), then $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has central vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has central vanishing order at least 3 at $s=1/2.$
Let $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). For the diagonal cycle, if $\Delta_{3}]_{f_{1},f_{2},f_{3}}\neq 0$ in $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$, the generalized Gross–Kudla conjecture and the conjectural nondegeneracy of the height pairing imply that $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\neq 0$. Then the Beilinson–Bloch conjecture predicts that $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$ has dimension 1 and is thus generated by $\Delta_{3}]_{f_{1},f_{2},f_{3}}$. Moreover, by (6.1), $L^{\prime}(1/2,\pi_{3})\neq 0$ (assuming $\epsilon(1/2,\pi_{3})=-1$). Then by the general Gross–Zagier formula [11, 25] and 2.2 (which applies by 3.2 (2)), we propose the following conjecture.
Conjecture 6.3.
Assume $\pi_{1}\simeq\pi_{2}^{\vee}$, and $\pi_{3}$ is self-dual with $\epsilon(1/2,\pi_{3})=-1$. Let $(A,\iota)$ corresponding to $\pi_{3}$ under 3.1 (2). Then there exists a zero cycle $Z$ on $X_{U}$ such that $\Delta_{3}]_{f_{1},f_{2},f_{3}}$ is a multiple of $\Delta_{2}]_{f_{1},f_{2}}\times{\mathfrak{pr}}_{\iota,*}f_{3,*}[Z]$. Moreover, we can let $Z$ be a CM 0-cycle associated to some quadratic CM extension of $F$.
The relation between modified diagonal cycles and CM cycles was already explored in [2], as well as used [12, 15]. Our conjecture gives a complete explaination.
7. An intriguing hyperelliptic Shimura curve of Dembélé
Let $F={\mathbb{Q}}\left(\cos\frac{1}{16}\pi\right)$. In this section, we describe a hyperelliptic Shimura curve $X_{U}$ over $F$ found by Dembélé [4, 5.6]. We first clarify that in [4, 5.6], a Shimura curve is of the form $X_{V}$ with $V$ open compact (instead of compact-modulo-center as in our paper) in $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. And the hyperelliptic curve constructed in loc. cit. is an Atkin–Lehner quotient of such a Shimura curve, which is still a Shimura curve in our sense.
Let us be more explicit. Let $\tau$ be any of the 8 infinite place of $F$. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. To keep notations close to [4], let us use $D$ to denote the quaternion algebra over $F$ being divison exactly at $\mathfrak{p}$} and all infinite places of $F$ except $\tau$. Let $V=\prod\limits_{v\text{ finite}}V_{v}\subset D^{\times}({\mathbb{A}}_{\mathrm{f}})$ be a maximal open compact subgroup. In other words, its $\mathfrak{p}$}-component is
$V_{\mathfrak{p}}=O_{{\mathfrak{p}}}^{\times},$
where $O_{\mathfrak{p}}$ is a maximal order of $D_{\mathfrak{p}}$, and
$V_{v}={\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ for }v\neq{\mathfrak{p}}.$
Since the narrow class number of $F$ is 1, $X_{V}$ is geometrically connected and $X_{V}\to X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ is an isomorphism. Here ${\mathbb{A}}_{\mathrm{f}}^{\times}$ is the center of $D^{\times}({\mathbb{A}}_{\mathrm{f}})$. This is the Shimura curve $X_{0}^{D}(1)$ in [4, Introduction, 4.1]. The generator of $D_{\mathfrak{p}}^{\times}/F_{\mathfrak{p}}^{\times}O_{{\mathfrak{p}}}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ normalizes $V_{\mathfrak{p}}F_{\mathfrak{p}}^{\times}$ and thus defines an automorphism of $X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ of order 2. (For example we can take the generator to be the uniformizer of a ramified quadratic extension of $F_{\mathfrak{p}}$ in $D_{\mathfrak{p}}$.) This is the Atkin–Lehner involution $w_{D}$ in [4, Introduction, 5.4]. In particular, the Atkin–Lehner quotient $X_{0}^{D}(1)/\langle{w_{D}}\rangle$ in [4, Introduction, 5.4] is $X_{U}$ with
$U_{\mathfrak{p}}=PD_{{\mathfrak{p}}}^{\times}$
and
$U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ for }v\neq{\mathfrak{p}}.$
Theorem 7.1 ([4, Theorem 5.9]).
The curve $X_{U}$ is hyperelliptic over F.
We only need the hyperellipticity over $\mathbb{C}$}.
By [4, 5.3], there are 40 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2, and they form 5 Galois orbits of sizes 4, 4, 4, 4 and 24 respectively. Moreover, let $w$ be the Atkin–Lehner involution acting on the $\Gamma_{0}({\mathfrak{p}})$-Hilbert modular forms. Then the newforms in the four Galois orbits of sizes 4 are exactly the ones with $w$-eigenvalue 1-1. Their Jacquet–Langlands correspondences to $D^{\times}$ exactly the ones with $w_{D}$-eigenvalue $1$. So these Jacquet–Langlands correspondences descend to $X_{U,{\mathbb{C}}}$ and span the space of holomorphic differential 1-forms on $X_{U,{\mathbb{C}}}$, which has genus 16. Then by 5.6, we have the following.
Theorem 7.2.
Consider the 16 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2 with $w$-eigenvalue 1-1. Let $f_{1},f_{2},f_{3}$ be any three of them. Then
${\mathrm{ord}}_{s=2}L\left(s,f_{1}\otimes f_{2}\otimes f_{3}\right)\geq 3.$
Remark 7.3.
(1) Computationally, it is not easy to verify whether the 3rd central derivatives of these $L$-functions vanish due to their large conductors.
(2) 6.2 suggests that we also should check whether the central derivatives of the $L$-functions vanish for single newforms in the theorem. This is easy for people with access to enough computational power.