### v2

# Hyperelliptic Shimura curves and $L$-functions of central vanishing order at least 3

Abstract.
We use hyperelliptic Shimura curves to find triple product $L$-functions of Hilbert newforms with central vanishing orders proved to be at least 3.
1. Introduction
In [10, 26], Gross and Zagier proved that the $L$-function of a certain modular elliptic curve over $\mathbb{Q}$} has central vanishing order 3. This was essential in Goldfeld’s effective solution to the Gauss class number problem for imaginary quadratic fields [5]. According to folklore conjectures, such $L$-functions with higher central vanishing orders are very rare. Moreover, such vanishing can not be proved by numerically computation. Thus, Gross and Zagier’s example is remarkable. In this paper, we present some analogs of Gross and Zagier’s example in triple product $L$-functions of abelian varieties.
Theorem 1.1.
Let $F={\mathbb{Q}}(\zeta_{32}+\zeta_{32}^{-1})$, the maximal totally real subfield of the cyclotomic field of 32nd roots of unity. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. Let $A_{1},A_{2},A_{3}$ be pairwise non-isomorphic 4-dimensional simple modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ and of conductor $\mathfrak{p}$}. Then
${\mathrm{ord}}_{s=2}L\left(s,h^{1}(A_{1})\otimes h^{1}(A_{2})\otimes h^{1}(A_{3})\right)\geq 3\cdot 4^{3}.$
Indeed, there are exactly four modular abelian varieties of ${\mathrm{GL}}_{2}$-type over $F$ of conductor $\mathfrak{p}$} and dimension 4, which are all simple. They exactly correspond to the only 4 Galois orbits of Hilbert newforms over $F$ with Atkin–Lehner sign 1-1. And each Galois orbit is of size 4. Then the fact that the triple product $L$-function has a holomorphic continuation to $s\in{\mathbb{C}}$ follows from [4, 16].
We actually prove a slightly more general version of 1.1 in terms of Hilbert newforms over $F$. In 8.2, we show that for any 3 pairwise distinct newforms of the 16 Hilbert newforms (in the four Galois orbits of size 4) above, their triple product $L$-function has central vanishing order at least 3.
In general, when some of the 3 newforms are the same, the triple product $L$-function is decomposable. See Section 6 . In this case, using usual modular forms over $\mathbb{Q}$}, we get many degree 6 $L$-functions with central vanishing order at least 2 (7.2).
Our proof of 1.1 is similar to the one of Gross and Zagier in spirit. They used their celebrated formula on modular curves, which relates $L$-functions of modular forms to the heights of (Heegner) CM 0-cycles. We use an analog proved by Yuan, S. Zhang, and W. Zhang [24] for some Shimura curves, which relates triples product $L$-functions to the heights of modified diagonal cycles on the triple product of Shimura curves. Both of us show that our cycles accidentally vanish for some reasons.
In our case, our cycles vanish due to an exceptional hyperelliptic involution, which is the work of Dembélé [3]. See Section 8. Here, an involution is exceptional if it does not come from an Atkin–Lehner involution. (In general, if an Atkin–Lehner involution is a hyperelliptic involution, it is not useful to produce higher central vanishing of $L$-functions. See 5.8.) Interestingly, the specific modular curve used for Gross and Zagier’s order 3 example, which is $X_{0}(37)$, is of genus 2. And its hyperelliptic involution is exceptional. It might be a coincidence, since the hyperellipticity is essential for our result, but it is not used in [10, 26].
Finally, we want to remark that there are only finitely many hyperelliptic Shimura curve, as shown in [19], let along the ones satisfying our specific criterions Section 5. Moreover, we are mostly interested in the ones of genus at least 3 due to the decomposability of triple product $L$-functions mentioned above. They are even rarer. For example, there are no such modular curves as we will see in Section 7. Such rareness is also a facet of the expectation that $L$-functions of central vanishing order at least 3 are very rare.
Structure of the paper
In Section 2 and Section 3, we review some background on abelian varieties and Shimura curves, respectively. In Section 4, we relate modified diagonals to the hyperellipticity of Shimura curves. In Section 5, we give criteria for triple product $L$-functions to have central vanishing order at least 3. In Section 6, we consider the cases where the triple product $L$-function is decomposable. In Section 7, we study examples of hyperelliptic modular curves and obtain many degree-6 $L$-functions over $\mathbb{Q}$} with central vanishing order at least 2. Finally, in Section 8, we study the hyperelliptic Shimura curve of Dembélé and prove 1.1.
Acknowledgments
The author would like to thank Peter Sarnak for raising the question of possible analogs of Gross–Zagier’s example in higher dimensions during a talk by the author at Princeton. The author is grateful to Yifeng Liu and Wei Zhang for many helpful discussions in the early stages of this work. The author also thanks Benedict Gross and Adam Logan for valuable communications that helped correct a mistake in the first version of this paper.
2. Triple product of abelian motives
For a smooth projective variety $V$ over a field, let ${\mathrm{Ch}}_{i}(V)$ be the Chow group of $V$ of dimension $i$ cycles with $\mathbb{Q}$}-coefficients, modulo rational equivalence. For a morphism $f:V\to U$, let $\Gamma_{f}\in{\mathrm{Ch}}_{*}(V\times U)$ be the graph of $f$.
Let $A$ be an abelian variety over a field $F$ of dimension $g$, with dual abelian variety $A^{\vee}$. Let ${\mathfrak{pr}}\in{\mathrm{Ch}}_{g}(A\times A)$ be the (2g1)(2g-1)-th Deninger–Murre projector, that is denoted by $\pi_{2g-1}$ in [2, Theorem 3.1]. For a variety $X$ over $F$, the abelian group structure of $A$ induces an abelian group structure on ${\mathrm{Mor}}(X,A)$, the set of morphisms of $F$-varieties. Then by regarding $X\times A$ as an abelian scheme over $X$, [22, Corollary A.3.2] implies that the following map is a homomorphism of abelian groups:
${\mathrm{Mor}}(X,A)$ $\to{\mathrm{Ch}}_{\dim X}(X\times A),$
$f$ $\mapsto{{\mathfrak{pr}}}\circ[\Gamma_{f}].$
Denote the unique $\mathbb{C}$}-linear extension ${\mathrm{Mor}}(X,A)_{\mathbb{C}}\to{\mathrm{Ch}}_{\dim X}(X\times A)$ by
$f\mapsto{{\mathfrak{pr}}_{f}}.$
For an idempotent $\iota\in{\mathrm{End}}(A)_{\mathbb{C}}$, by [2, Proposition 3.3], ${\mathfrak{pr}}_{\iota}$ is a projector. Moreover, if $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$, one can easily check that its image $\iota^{\dagger}$ under a Rosati involution on ${\mathrm{End}}(A)_{\mathbb{C}}$ associated with a polarization does not depend on the choice of the polarization.
Proposition 2.1.
Let the symmetric group $S_{n}$ act on $A^{n}$ by permuting the components. Assume that $\dim\iota_{*}H^{1}(A)<n$. For $z\in{\mathrm{Ch}}_{*}(A^{n})$ invariant by $S_{n}$, we have $({\mathfrak{pr}}_{\iota}^{n})_{*}z=0$.
Proof.
The proof is the same as [20, Corollary 2.2.3]. ∎
Proposition 2.2.
Let $f:X\to A$ be a morphism from a smooth projective curve over $F$ whose associated Albanese morphism to $A$ is surjective. Let $\Delta\subset X^{2}$ be the diagonal. Assume that $\iota$ is in the center of ${\mathrm{End}}(A)_{\mathbb{C}}$ such that $\iota^{\dagger}=\iota$ and $\dim\iota_{*}H^{1}(A)\leq 2$. Then $({\mathfrak{pr}}_{\iota}^{2})_{*}{\mathrm{Ch}}_{*}(A^{2})={\mathbb{C}}\cdot({\mathfrak{pr}}_{\iota}^{2})_{*}(f^{2})_{*}[\Delta]$, which is nonzero.
Proof.
The proof is the same as [20, Corollary 2.2.7 (1)]. ∎
Remark 2.3.
The relation between the above two propositions and [20, Corollary 2.2.3 , Corollary 2.2.7 (1)] is not just similar. Indeed, if $A$ is simple, we have an isomoprhism between the motives $(A,{\mathfrak{pr}}_{\iota})$ and $(X,\delta_{f})$ defined in [20, Section 2.4][1][1][1]In loc. cit., we used $\phi$ instead of $f$ to denote a morphism from $X$ to $A$. . This isomoprhism can be easily proved using the decomposition of $\delta_{f}$ in (the summand on the right hand side of) [22, (2.13)] and the nilpotency of Kimura [13, Proposition 7.5].
3. Shimura curves
Let $F$ be a totally real number field and ${\mathbb{A}}_{\mathrm{f}}$ its ring of finite adeles. Let $B$ be a quaternion algebra split at one infinite place $\tau:F\hookrightarrow{\mathbb{R}}$ of $F$ and division at all other infinite places of $F$. Let $B^{+}\subset B^{\times}$ be the subgroup of elements with totally positive norms, equivalently, with positive determinants in $B_{\tau}^{\times}\simeq{\mathrm{GL}}_{2}({\mathbb{R}})$. Let $\mathbb{H}$} be the complex upper half-plane, and let $B^{+}\subset{\mathrm{GL}}_{2}({\mathbb{R}})$ act on $\mathbb{H}$} by the fractional linear transformation.
For an open compact-modulo-center subgroup $U$ of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, we have the smooth compactified Shimura curve $X_{U}$ for $B^{\times}$ of level $U$ over $F$ [25, 3.1.4][2][2][2]Technically, sometimes by “Shimura curve $X_{U}$ for $B^{\times}$”, it is required that $U$ is compact. And if $U$ contains the center of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, $X_{U}$ should be refered to a Shimura curve for $PB^{\times}$. We do not make this distinction here.. The complex uniformization of $X_{U}$ via the distinguished infinite place $\tau:F\hookrightarrow{\mathbb{R}}\subset{\mathbb{C}}$ is given by
(3.1) $X_{U,\tau,{\mathbb{C}}}\simeq B^{+}\backslash{\mathbb{H}}\times B^{\times}({\mathbb{A}}_{\mathrm{f}})/U\coprod\{\text{cusps}\}.$
Here, the cusps exist if and only if $X_{U}$ is a modular curve, i.e. $F={\mathbb{Q}}$, and $B$ is the matrix algebra.
[…]
For $x\in X_{U}({\mathbb{C}})$ that is the image of a point in ${\mathbb{H}}\times G({\mathbb{A}}_{\mathrm{f}})/U$ with a nontrivial stabilizer in $B^{+}$, we call $x$ an elliptic point and let $r_{x}$ be the order of the stabilizer. There are only finitely many elliptic points on $X_{U}$. Let
${\mathfrak{e}}_{U}=\sum_{\text{elliptic }x}\left(1-\frac{1}{r_{x}}\right)x$
Let $K_{U}$ be a canonical divisor of $X_{U}$, and ${\mathcal{K}}_{U}=K_{U}+{\mathfrak{e}}_{U}$ an orbifold canonical divisor of $X_{U}$. Let ${\mathfrak{c}}_{U}$ be the divisor of the cusps of $X_{U}$, which is 0 unless $X_{U}$ is a modular curve. Let
$L_{U}={\mathcal{K}}_{U}+{\mathfrak{c}}_{U}.$
[…]
$\lambda_{U}=[L_{U}]/\deg L_{U}^{\circ}$
where $L_{U}^{\circ}$ is the restriction of $L_{U}$ to any geometrically connected component of $X_{U}$. In [22, Remark 3.2.4], we further noticed that if $X_{U}$ is a modular curve, then ${\mathfrak{c}}_{U}$ is a multiple of $\lambda_{U}$. So
(3.2) $\lambda_{U}=[{\mathcal{K}}_{U}]/\deg{\mathcal{K}}_{U}^{\circ},$
provided that $\deg{\mathcal{K}}_{U}>0$.
For an abelian variety $A$ over $F$, let ${\mathrm{Mor}}^{0}(X_{U},A)$ be the subspace of $f\in{\mathrm{Hom}}(X_{U},A)_{\mathbb{Q}}$ such that for some (equivalently for every) representative $\sum_{i}a_{i}p_{i}$ of $\lambda_{U,\overline{F}}$, where $a_{i}\in{\mathbb{Q}}$ and $p_{i}\in X_{U}(\overline{F})$, the sum of $f(p_{i})\otimes a_{i}$ over $i$’s in the group $A(\overline{k})\otimes{{\mathbb{Q}}}$ is the identity of $A$. Let
[…]
$\pi_{A}^{U}={\mathrm{Mor}}^{0}(X_{U},A).$
Let ${\mathrm{End}}(A)_{\mathbb{Q}}$ act on $\pi_{A}^{U}$ and $\pi_{A}$ from left by post-composition, and one may regard $\pi_{A}$ as a $B^{\times}({\mathbb{A}}_{\mathrm{f}})$-representation with ${\mathrm{End}}(A)_{\mathbb{Q}}$-coefficients.
Theorem 3.1 ([25, 3.2.2, 3.2.3]).
Let $A$ be a simple abelian variety over $F$ such that $\pi_{A}\neq 0$.
(1) The $\mathbb{Q}$}-algebra ${\mathrm{End}}(A)_{\mathbb{Q}}$ is a finite field extension of $\mathbb{Q}$} and ${\mathrm{End}}(A)_{\mathbb{Q}}:{\mathbb{Q}}]=\dim A$.
(2) Consider the set of pairs $(A,\iota)$ where $A$ is a simple abelian variety over $F$ such that $\pi_{A}\neq 0$ and $\iota:{\mathrm{End}}(A)_{\mathbb{Q}}\hookrightarrow{\mathbb{C}}$ is an embedding. The map $(A,\iota)\to\pi_{A,\iota}$ is a bijection to the set of finite components of automorphic representations of $B^{\times}$, whose Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is cuspidal and holomorphic of weight 2.
Remark 3.2.
(1) By (1) of the theorem, $A$ is of ${\mathrm{GL}}_{2}$-type. A simple abelian variety over $F$ of ${\mathrm{GL}}_{2}$-type with $\pi_{A}\neq 0$ for some $B$ as above is called modular (or automorphic in [25]).
(2) Since ${\mathrm{End}}(A)_{\mathbb{Q}}$ is a field so that irreducible idempontents of ${\mathrm{End}}(A)_{\mathbb{Q}}$ are in bijection with embeddings $M_{A}\hookrightarrow{\mathbb{C}}$. Moreover, they are automatically central. Thus, here we use the same notation $\iota$ as in 2.1 and $2.$
Below we simplify notations and let $\lambda,{\mathfrak{c}},{\mathfrak{e}}$ denote $\lambda_{U},{\mathfrak{c}}_{U},{\mathfrak{e}}_{U}$ if $U$ is clear from the context. Assume $\deg(K_{U})>0$. Let
$\xi=[K_{U}]/\deg K_{U}^{\circ}.$
By definition, ${\mathfrak{pr}}_{f,*}\lambda=0$. Then by (3.2), we have
(3.3) ${\mathfrak{pr}}_{f,*}\xi=0\text{ if and only if }{\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0.$
4. Diagonals and hyperellipticity
Consider the $n$-th power of the Shimura curve $X_{U}$. Let
[…]
where the transition maps in the inverse limit are pushforwrds by the natural morphisms between Shimura curves. Let $B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ act on $\overline{{\mathrm{Ch}}}(n)$ (from left) by pullback. Let $\pi_{i}$ be of the form $\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2), which is a representation of $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. We have a linear map $\pi_{i}^{U}\to{\mathrm{Ch}}_{1}(X_{U}\times A_{i})$, sending $f_{i}$ to ${\mathfrak{pr}}_{f_{i}}$. Let
$\boxtimes_{i=1}^{n}\pi_{i}^{U}$ $\to\otimes_{i=1}^{n}{\mathrm{Ch}}_{1}(X_{U}\times A_{i}),$
$f$ $\mapsto{\mathfrak{pr}}_{f}.$
be the tensor product of these maps.
The following is a slight generalization of [22, Theorem 3.3.2]. The proof is the same.
Theorem 4.1.
Let $H\subset B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$ be a subgroup. Let $\Pi\subset\boxtimes_{i=1}^{n}\pi_{i}$ be a subspace be stable by $H$. If $\Pi$ has no nonzero $H$-invariant linear forms, then for $z=(z_{U})_{U}\in{\mathrm{Ch}}(n)$ invariant by $H$, we have
${\mathfrak{pr}}_{f,*}z=0$
for all $f\in\Pi$.
In particular, the theorem applies to a sub-$B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$-representation $\Pi$ of $\boxtimes_{i=1}^{n}\pi_{i}$.
Let $\Delta_{n}=\Delta_{U,n}$ be the diagonal $X_{U}^{n}$. Let
$\Delta_{n}]_{f}={\mathfrak{pr}}_{f,*}[\Delta_{n}]$
in ${\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times...\times A_{n}).$
Accordingly, we have the following slight generalization of [22, Corollary 3.3.4].
Corollary 4.2.
If a sub-$B^{\times,n}({\mathbb{A}}_{\mathrm{f}})$-representation $\Pi$ of $\boxtimes_{i=1}^{n}\pi_{i}$. has no nonzero diagonal-$B^{\times}({\mathbb{A}}_{\mathrm{f}})$-invariant linear forms, then
$\Delta_{n}]_{f}=0$
for all $f\in\Pi$.
Remark 4.3.
In the case $n=3$ and $\Pi=\boxtimes_{i=1}^{n}\pi_{i}$, this condition has been extensively studied in [19].
Here are some easy examples for applying this corollary to $\Pi=\boxtimes_{i=1}^{n}\pi_{i}$.
Example 4.4.
(1) If $n=1$, let $H=B^{\times}({\mathbb{A}}_{\mathrm{f}})$ and we have $\Delta_{1}]_{f_{1}}=0$.
(2) If $n=2$, $H$ is the diagonal $B^{\times}({\mathbb{A}}_{\mathrm{f}})$, and $z_{U}=[\Delta_{2}]$, then $\Delta_{2}]_{f_{1}\otimes f_{2}}=0$ if $\pi_{1}\not\simeq\pi_{2}^{\vee}$.
(3) If the product of the central characters of $\pi_{i}$’s is not trivial, then $\Delta_{n}]_{f_{1}\otimes f_{2},...\otimes f_{n}}=0$.
For a subset $I$ of $\{1,2,...,n\}$, let $p_{I}$ be the projection of $X_{U}^{n}$ to the product of the components indexed by $I$.
Proposition 4.5.
(1) Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then
$\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ I=\{a,b\},J=\{c\}\end{subarray}}p_{I}^{*}[\Delta_{2}]_{f_{a}\otimes f_{b}}\cdot p_{J}^{*}{\mathfrak{pr}}_{f_{c},*}\xi$
for all $f_{i}\in\pi_{i}^{U}$.
(2) Further assume $\pi_{3}\not\simeq\pi_{1}^{\vee}$ and $\pi_{3}\not\simeq\pi_{2}^{\vee}$. Then
$\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=-p_{1,2}^{*}[\Delta_{2}]_{f_{1}\otimes f_{2}}\cdot p_{3}^{*}{\mathfrak{pr}}_{f_{3},*}\xi,$
for all $f_{i}\in\pi_{i}^{U}$.
(3) Further assume $\pi_{i}\not\simeq\pi_{j}^{\vee}$ for any $i\neq j$ in $\{1,2,3\}$. Then
$\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$
for all $f_{i}\in\pi_{i}^{U}$.
Proof.
The modified diagonal defined by Gross and Schoen [9, 27] is the algebraic cycle
(4.1) $[\Delta_{3}]_{\xi}=[\Delta_{3}]-\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ |I|=2\end{subarray}}p_{I}^{*}[\Delta_{2}]\cdot p_{J}^{*}\xi+\sum_{\begin{subarray}{c}I\coprod J=\{1,2,3\}\\ |I|=1\end{subarray}}p_{I}^{*}[\Delta_{1}]\cdot p_{J}^{*}(\xi\times\xi)$
By hyperellipticity and [9, 4.5], $\Delta_{\xi}]=0$. The proposition follows from 4.4 (1)(2). ∎
Combining (1) and (2) of the proposition with (3.3), we have a corollary.
Corollary 4.6.
(1) Assume every geometrically connected component of $X_{U}$ is hyperelliptic. Then $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$ for $f_{i}\in\pi_{i}^{U}$ such that ${\mathfrak{pr}}_{f_{i},*}[{\mathfrak{e}}]=0$.
(2) Further assume $\pi_{3}\not\simeq\pi_{1}^{\vee}$ and $\pi_{3}\not\simeq\pi_{2}^{\vee}$. Then $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$ for $f_{i}\in\pi_{i}^{U}$ such that ${\mathfrak{pr}}_{f_{3},*}[{\mathfrak{e}}]=0.$
Remark 4.7.
Note that elliptic points are CM points and $\mathfrak{e}$} is a sum of CM cycles. Then the equation ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$, if holds, can usually be verified by local conditions. See Section 7.
5. Triple product $L$-functions of central vanishing order at least 3
We keep the notations from the last section, except that from now on, the notation $\pi$ or $\pi_{i}$ will denote an automorphic representation of $B^{\times}$, instead of its finite part. We always assume that it is cuspidal holomorphic of weight $2$ in the sense that its Jacquet–Langlands correspondence to ${\mathrm{GL}}_{2,F}$ is a cuspidal automorphic representation holomorphic of weight $2$.
For a place $v$ of $F$, let $\epsilon(B_{v})=1$ if is $B_{v}$ is split, and let $\epsilon(B_{v})=-1$ if is $B_{v}$ is division. So if $v=\tau$, $\epsilon(B_{v})=1$ and for any othet infinite place $v$, $\epsilon(B_{v})=-1.$ Then
$\prod_{v}\epsilon(B_{v})=1.$
Here the product is over all places of $F$. Let $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})$ be the local root number of $\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v}$. A result of Prasad [17, 18] states the following.
Theorem 5.1.
If
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\},$
then it has dimension 1 and $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=\epsilon(B_{v})$.
Corollary 5.2.
If
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\}$
for all finite places $v$ of $F$, then $\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Proof.
At every infinite place $v$, $\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-1$ (see [17, Section 9]). At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. So by 5.1, we have
$\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=\prod_{v}\epsilon(1/2,\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v})=-\prod_{v}\epsilon(B_{v})=-1.$
Let $S_{B}$ be the finite set of finite places $v$ of $F$ such that $B_{v}$ is a division quaternion algebra. Below we only consider a maximal open compact subgroup $U=\prod\limits_{v\text{ finite}}U_{v}\subset PB^{\times}({\mathbb{A}}_{\mathrm{f}})$ until the final remark (5.8). In other words, its $v$-component is
(5.1) $U_{v}=PB_{v}^{\times}\text{ for }v\in S_{B},$
and
(5.2) $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ or }F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right),g_{v}}\rangle\text{ for }v\not\in S_{B},$
where ${\mathfrak{m}}_{F_{v}}$ is the maximal ideal of ${\mathcal{O}}_{F_{v}}$, $\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right)$ is the subgroup of matrices in ${\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ that are upper triangular modulo ${\mathfrak{m}}_{F_{v}}$ and $g_{v}\in\begin{bmatrix}0&1\\ {\mathcal{O}}_{F_{v}}\backslash{\mathfrak{m}}_{F_{v}}&0\end{bmatrix}$. If $\pi^{U}\neq 0$, then $\pi$ has trivial central character and thus self-dual, and $\dim\pi_{v}^{U_{v}}=1$ for $v$ finite. Explicitly, for $v\in S_{B}$, $\pi_{v}$ is the trivial representation of $B_{v}^{\times}$. Its Jacquet-Langlands correspondence to ${\mathrm{GL}}_{2}(F_{v})$ is the the Steinberg representation. In this cases, the root number $\epsilon(1/2,\pi_{v})=-1$. For $v\not\in S_{B}$, if $U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)$ for $v\not\in S_{B}$, then $\pi_{v}$ is an unramified representation of ${\mathrm{GL}}_{2}(F_{v})$. If $U_{v}=F_{v}^{\times}\langle{\Gamma_{0}\left({\mathfrak{m}}_{v}\right),g_{v}}\rangle$, then $\pi_{v}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of $F_{v}^{\times}$. In both cases, the root number $\epsilon(1/2,\pi_{v})=1$.
This lemma is not needed for the results in this section. We will use it later.
Lemma 5.3.
If $\pi^{U}\neq 0$, then $\epsilon(1/2,\pi)=-1$.
Proof.
At an infinite place, $\epsilon(1/2,\pi_{v})=-1$. At every infinite place $v$, $\epsilon(B_{v})=-1$ for except one place. And at every infinite place $v$, as we have seen above the lemma, $\epsilon(1/2,\pi_{v})=\epsilon(B_{v})$. So we have the first of the following equations $\epsilon(1/2,\pi)=-\prod_{v}\epsilon(B_{v})=-1.$
This lemma is essential for the results in this section.
Lemma 5.4.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$ and $v$ finite, then
${\mathrm{Hom}}_{B_{v}^{\times}}(\pi_{1,v}\otimes\pi_{2,v}\otimes\pi_{3,v},{\mathbb{C}})\neq\{0\}.$
For $\ell\neq 0$ in it and $\phi_{i}\neq 0\in\pi_{i,v}^{U_{v}}$ (both unique up to scalar as seen above), we have
(5.3) $\ell(\phi_{1}\otimes\phi_{2}\otimes\phi_{3})\neq 0.$
Proof.
If $v\in S_{B}$ so that $\pi_{v}=1$, the lemma is trivial. Otherwise, the lemma is [8, Proposition 6.1, 6.3]. ∎
Then by 5.2, we have the following
Corollary 5.5.
If $\pi_{i}^{U}\neq 0$ for $i=1,2,3$, then $\epsilon(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=-1$.
Let $\pi_{i,{\mathrm{f}}}$ denote the finite component of $\pi_{i}$. Then $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2).
The main theorem of this section is the following.
Theorem 5.6.
Assume
  • $U$ satisfies (5.1) and (5.2),
  • every geometrically connected component of $X_{U}$ is hyperelliptic, and
  • $\pi_{i}^{U}\neq 0$ for $i=1,2,3$.
If $\pi_{i}$’s are pairwise non-isomorphic, then
${\mathrm{ord}}_{s=1/2}L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\geq 3.$
Proof.
The generalized Gross–Kudla conjecture [6] is proved in [24] under the first two assumptions in the theorem. By this and (5.3), the central derivative $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ is a multiple of the height of $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$, where $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, unique up to scalar. But by 4.5 (3), $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$. Now the theorem follows from $5.$ ∎
We will use this theorem to prove 1.1.
Replacing 4.5 (3) by 4.6 (2), we have the following more general theorem.
Theorem 5.7.
Assume
  • $U$ satisfies (5.1) and (5.2),
  • every geometrically connected component of $X_{U}$ is hyperelliptic, and
  • $\pi_{i}^{U}\neq 0$ for $i=1,2,3$,
  • $\pi_{3}\not\simeq\pi_{1}$, $\pi_{3}\not\simeq\pi_{2}$.
Let $f_{3}\neq 0\in\pi_{3,{\mathrm{f}}}^{U}$, unique up to scalar. If ${\mathfrak{pr}}_{f_{3},*}[{\mathfrak{e}}]=0$, then
${\mathrm{ord}}_{s=1/2}L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\geq 3.$
When $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}$, $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has central vanishing order at least 3 automatically. We discuss this in the coming section.
Remark 5.8.
(1) The condition we impose on the levels of $\pi_{i}$’s, i.e., on $U$, above is for the purpose of getting 5.4, especially (5.3). Slightly more generally, we may further allow $U_{v}$ to be $F_{v}^{\times}{\mathcal{O}}_{B_{v}}^{\times}$ for $v\in S_{B}$ and the maximal order ${\mathcal{O}}_{v}$ of $B_{v}$, and allow $U_{v}$ to be $F_{v}^{\times}\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}\right)$ for $v\not\in S_{B}$. Then we need to match the representations so that the lemma still holds. For example, for $v\in S_{B}$, let $\pm 1$ be the two representations of $B_{v}^{\times}/F_{v}^{\times}O_{v}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$. Then we need the product of $\pi_{i,v}$’s to be 1. In particular, if the Atkin–Lehner involution, defined by the generator $B_{v}^{\times}/F_{v}^{\times}O_{v}^{\times}$, is hyperelliptic, then all $\pi_{i,v}$’s are 1-1. And the lemma will not hold.
(2) For $v\not\in S_{B}$, more generally, the Atkin–Lehner involution is defined if $U_{v}=F_{v}^{\times}\Gamma_{0}\left({\mathfrak{m}}_{F_{v}}^{n_{v}}\right)$ for some positive integer $n_{v}$. Indeed, any $g_{v}\in\begin{bmatrix}0&1\\ {\mathfrak{m}}_{F_{v}}^{n-1}\backslash{\mathfrak{m}}_{F_{v}}^{n}&0\end{bmatrix}$ normalizes $U_{v}$ and defines the same Atkin–Lehner involution $w_{v}$ on $X_{U}$ at $v$. If $\dim\pi_{i,v}^{U_{v}}=1$, then $w_{v}$ acts on $\pi_{i,v}^{U_{v}}$ by $\pm 1$. If the action is by 1-1 and $f_{i}\neq 0\in\pi_{i,{\mathrm{f}}}^{U}$, one can use $w_{v}$ to show that $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}=0$. Indeed, $(w_{v},w_{v},w_{v})$ fixes $\Delta_{3}]$ and thus fixes $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$, but sends each $f_{i}$ to $f_{i}$. The case when $w_{v}$ is a hyperelliptic involution is simply the case that any $\pi$ with $\pi_{i,{\mathrm{f}}}^{U}\neq 0$ satisfies that $\dim\pi_{i,v}^{U_{v}}=1$ and $w_{v}$ acts on $\pi_{i,v}^{U_{v}}$ by 1-1. However, in this case, (5.3) fails by the same argument. Thus $w_{v}$ being a hyperelliptic involution can not be used to produce higher central vanishing of $L$-functions.
(3) Finally, the above remarks on an Atkin–Lehner involution at a single finite place extend to a product of Atkin–Lehner involutions at several finite places in an obvious way.
6. Decomposable triple product $L$-functions
With the generalized Gross–Kudla conjecture[6, 24], we have used the (modified) diagonal cycle to study triple product $L$-functions. Now we take the opposite direction, and want to understand the behavior of the diagonal cycle for a general Shimura curve predicted by triple product $L$-functions. Keep the notations from the last section, except that now we abandon (5.1) and (5.2). We analyze two cases when the triple product $L$-function is decomposable. Such consideration appeared in an earlier version of [24] for the heights of the diagonal cycle.
First, if $\pi_{1}\simeq\pi_{2}^{\vee}=:\pi$, then
(6.1) $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=L(s,{\mathrm{ad}}\pi\otimes\pi_{3})L(s,\pi_{3}).$
Assume $\pi_{3}$ is self-dual with root number $\epsilon(1/2,\pi_{3})=-1$. Then $L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})$ has vanishing order at least 3 at $s=1/2$ if and only if $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has vanishing order at least 2 at $s=1/2$ or $L(s,\pi_{3})$ has vanishing order at least 3 at $s=1/2.$ Then by 5.3, we have the following.
Proposition 6.1.
In the situation of 5.7, if $\pi_{1}\simeq\pi_{2}=:\pi$ (which are self-dual) and $L^{\prime}(1/2,\pi_{3})\neq 0$, then $L(s,{\mathrm{ad}}\pi\otimes\pi_{3})$ has vanishing order at least 2 at $s=1/2$.
Now, we consider the diagonal cycle. Let $\pi_{i,{\mathrm{f}}}=\pi_{A_{i},\iota_{i}}$ for some pair $(A_{i},\iota_{i})$ as in 3.1 (2). If $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}\neq 0$ in $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$, the generalized Gross–Kudla conjecture and the conjectural nondegeneracy of the height pairing imply that $L^{\prime}(1/2,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})\neq 0$. Then the Beilinson–Bloch conjecture predicts that $(\iota_{1}\times\iota_{2}\times\iota_{3})_{*}{\mathrm{Ch}}_{1}(A_{1}\times A_{2}\times A_{3})$ has dimension 1 and is thus generated by $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$. Moreover, by (6.1), $L^{\prime}(1/2,\pi_{3})\neq 0$ (assuming $\epsilon(1/2,\pi_{3})=-1$). Then by the general Gross–Zagier formula [10, 25] and 2.2 (which applies by 3.2 (2)), we propose the following conjecture.
Conjecture 6.2.
Assume $\pi_{1}\simeq\pi_{2}^{\vee}$, and $\pi_{3}$ is self-dual with $\epsilon(1/2,\pi_{3})=-1$. Let $(A,\iota)$ corresponding to $\pi_{3}$ under 3.1 (2). Then there exists a zero cycle $Z$ on $X_{U}$ such that $\Delta_{3}]_{f_{1}\otimes f_{2}\otimes f_{3}}$ is a multiple of $\Delta_{2}]_{f_{1}\otimes f_{2}}\times{\mathfrak{pr}}_{\iota,*}f_{3,*}[Z]$. Moreover, we can let $Z$ be a CM 0-cycle associated to some quadratic CM extension of $F$.
The relation between modified diagonal cycles and CM cycles was already explored in [1], as well as used [12, 15]. Our conjecture gives a complete explanation. Alternatively, in an earlier version of [24], the authors proposed to use an ample and symmetric line bundle associated to the polarization $A_{1}\to A_{2}\simeq A_{1}^{\vee}$ to construct 0-cycles that replace CM cycles in this case. See also [28, 5.3].
Second, if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual, then
$L(s,\pi_{1}\otimes\pi_{2}\otimes\pi_{3})=L(s,{\mathrm{sym}}^{3}\pi)L(s,\pi)^{2}.$
So if the root number $\epsilon(1/2,\pi)=-1$, then $L(s,\pi^{\otimes 3})$ has vanishing order at least 2 at $s=1/2$. Thus by 5.3 and 5.5, 5.6 is trivial if $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}.$
For the diagonal cycle, motivated by the generalized Gross–Kudla conjecture, we propose the following conjecture.
Conjecture 6.3.
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$ is self-dual and $\epsilon(1/2,\pi)=-1$, then $\Delta_{3}]_{f}=0$ for all $f\in\boxtimes_{i=1}^{3}\pi_{i}^{U}$.
2.1 (which applies by 3.2 (2)) gives the following proposition. It contains a special case of 6.3.
Proposition 6.4.
If $\pi_{1}\simeq\pi_{2}\simeq\pi_{3}=:\pi$, for $f\in{\mathrm{sym}}^{3}\pi$, $\Delta_{3}]_{f}=0$.
Remark 6.5.
If $\epsilon(1/2,\pi)=1$, the vanishing in the proposition follows from 4.2 and a result of Prasad [17, Theorem 6]. This is purely local and representation-theoretical.
7. Hyperelliptic modular curves
Consider the case of modular curves, i.e., let $B={\mathrm{M}}_{2}({\mathbb{Q}}).$ A Shimura curve $X_{U}$ with $U$ satisfying (5.2) is the quotient $X_{0}^{*}(N)$ of the classical modular curve $X_{0}(N)$ by the full Atkin–Lehner group, for some $N$ square free. (See 5.8 (2) for an adelic definition of Atkin–Lehner involutions.) Let $U(N)\subset{\mathrm{PGL}}_{2}({\mathbb{A}}_{\mathrm{f}})$ be the corresponding open compact subgroup.
By [11], the list of square free $N$’s with $X_{0}^{*}(N)$ hyperelliptic consists of the following 39 numbers:
[…]
Moreover, these $X_{0}^{*}(N)$’s all have genus 2. Thus we can not apply 5.6. Instead, we want to apply 5.7. We need the following lemma.
Lemma 7.1.
Let $N$ be square free, Assume
  • (1)
    there exists a prime factor $p\mid N$ that is inert in $\mathbb{Q}(i)$, and
  • (2)
    there exists a (possibly equal) prime factor $q\mid N$ that is inert in $\mathbb{Q}(\sqrt{-3})$.
For a cuspidal automorphic representation $\pi$ of ${\mathrm{GL}}_{2,{\mathbb{Q}}}$ holomorphic of weight $2$, if $f\neq 0\in\pi^{U(N)}$, then ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}]=0$ on $X_{0}^{*}(N)$.
Proof.
The divisor of elliptic points has a decomposition ${\mathfrak{e}}={\mathfrak{e}}_{2}+{\mathfrak{e}}_{3}$ into a CM divisor with CM by $\mathbb{Q}(i)$ and $\mathbb{Q}(\sqrt{-3})$. We show ${\mathfrak{pr}}_{f,*}[{\mathfrak{e}}_{2}]={\mathfrak{pr}}_{f,*}[{\mathfrak{e}}_{3}]=0$. We prove the first equation and the proof of the second is the same.
Let $E=\mathbb{Q}(i)$ embedded in $B={\mathrm{M}}_{2}({\mathbb{Q}})$, uniquely up to conjugation. Using the formulation in [25, 1.3.2], we only need to show ${\mathrm{Hom}}_{E_{p}^{\times}}(\pi_{p},{\mathbb{C}})=0.$ Indeed, $\pi_{p}$ is the the Steinberg representation twisted by the unique nontrivial unramified quadratic character of ${\mathbb{Q}}_{p}^{\times}$. Let $D_{p}$ be the unique division quaternion algebra over ${\mathbb{Q}}_{p}$ and $E_{p}$ embedded in $D_{p}$, uniquely up to conjugation. The Jacquet–Langlands correspondence of $\pi_{p}$ to $D_{p}^{\times}$ is the unique nontrivial (quadratic) character $\chi$ of $D_{p}^{\times}/{\mathbb{Q}}_{p}^{\times}O_{p}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$, where $O_{p}$ is a maximal order of $D_{p}$. Since ${\mathrm{Hom}}_{E_{p}^{\times}}(\chi,{\mathbb{C}})\neq 0$ by assumption (1), by the Tunnell–Saito dicohotmy [25, Theorem 1.3], ${\mathrm{Hom}}_{E_{p}^{\times}}(\pi_{p},{\mathbb{C}})=0.$ The lemma is proved. ∎
In $\mathbb{Q}(i)$, a prime $p$ is inert if and only if $p\equiv 3\pmod{4}$. In $\mathbb{Q}(\sqrt{-3})$, a prime $q$ is inert if and only if $q\equiv 2\pmod{3}$. Among the $39$ squarefree $N$ in the list, the following $26$ contain at least one prime factor $p\equiv 3\pmod{4}$ and at least one prime factor $q\equiv 2\pmod{3}$.
Table 1. Squarefree NN in Equation 7.1 with prime factors inert in $\mathbb{Q}(i)$ and prime factors inert in $\mathbb{Q}(\sqrt{-3})$.
[…]
Theorem 7.2.
For $N$ in Table 1, let $f,g$ be (the only) two distinct newforms of level $\Gamma_{0}(N)$ and Atkin–Lehner sign +1+1 at all primes dividing $N$. Then
${\mathrm{ord}}_{s=2}L\left(s,{\mathrm{sym}}^{2}f\otimes g\right)\geq 2.$
Proof.
By 7.1, 5.7 applies. By 6.1, we only need to show that $L^{\prime}(1,g)\neq 0$. This can be read from LMFDB [14]. ∎
8. An intriguing hyperelliptic Shimura curve of Dembélé
Let $F={\mathbb{Q}}(\zeta_{32}+\zeta_{32}^{-1})$, the maximal totally real subfield of the cyclotomic field of 32nd roots of unity. In this section, we describe a hyperelliptic Shimura curve $X_{U}$ over $F$ found by Dembélé [3, 5.6]. We first clarify that in [3, 5.6], a Shimura curve is of the form $X_{V}$ with $V$ open compact (instead of compact-modulo-center as in our paper) in $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. And the hyperelliptic curve constructed in loc. cit. is an Atkin–Lehner quotient of such a Shimura curve, which is still a Shimura curve in our sense.
Let us be more explicit. Let $\tau$ be any of the 8 infinite place of $F$. Let $\mathfrak{p}$} be the unique prime of $F$ above 2. To keep notations close to [3], let us use $D$ to denote the quaternion algebra over $F$ being divison exactly at $\mathfrak{p}$} and all infinite places of $F$ except $\tau$. Let $V=\prod\limits_{v\text{ finite}}V_{v}\subset D^{\times}({\mathbb{A}}_{\mathrm{f}})$ be a maximal open compact subgroup. In other words, its $\mathfrak{p}$}-component is
$V_{\mathfrak{p}}=O_{{\mathfrak{p}}}^{\times},$
where $O_{\mathfrak{p}}$ is a maximal order of $D_{\mathfrak{p}}$, and
$V_{v}={\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ for }v\neq{\mathfrak{p}}.$
Since the narrow class number of $F$ is 1, $X_{V}$ is geometrically connected and $X_{V}\to X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ is an isomorphism. Here ${\mathbb{A}}_{\mathrm{f}}^{\times}$ is the center of $D^{\times}({\mathbb{A}}_{\mathrm{f}})$. This is the Shimura curve $X_{0}^{D}(1)$ in [3, Introduction, 4.1]. The generator of $D_{\mathfrak{p}}^{\times}/F_{\mathfrak{p}}^{\times}O_{{\mathfrak{p}}}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ normalizes $V_{\mathfrak{p}}F_{\mathfrak{p}}^{\times}$ and thus defines an automorphism of $X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ of order 2. (For example we can take the generator to be the uniformizer of a ramified quadratic extension of $F_{\mathfrak{p}}$ in $D_{\mathfrak{p}}$.) This is the Atkin–Lehner involution $w_{D}$ in [3, Introduction, 5.4]. In particular, the Atkin–Lehner quotient $X_{0}^{D}(1)/\langle{w_{D}}\rangle$ in [3, Introduction, 5.4] is $X_{U}$ with
$U_{\mathfrak{p}}=PD_{{\mathfrak{p}}}^{\times}$
and
$U_{v}=F_{v}^{\times}{\mathrm{GL}}_{2}\left({\mathcal{O}}_{F_{v}}\right)\text{ for }v\neq{\mathfrak{p}}.$
Theorem 8.1 ([3, Theorem 5.9]).
The curve $X_{U}$ is hyperelliptic over F.
We only need the hyperellipticity over $\mathbb{C}$}.
By [3, 5.3], there are 40 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2, and they form 5 Galois orbits of sizes 4, 4, 4, 4 and 24 respectively. Moreover, let $w$ be the Atkin–Lehner involution acting on the $\Gamma_{0}({\mathfrak{p}})$-Hilbert modular forms. Then the newforms in the four Galois orbits of sizes 4 are exactly the ones with $w$-eigenvalue 1-1. Their Jacquet–Langlands correspondences to $D^{\times}$ exactly the ones with $w_{D}$-eigenvalue $1$. So these Jacquet–Langlands correspondences descend to $X_{U,{\mathbb{C}}}$ and span the space of holomorphic differential 1-forms on $X_{U,{\mathbb{C}}}$, which has genus 16. Then by 5.6, we have the following.
Theorem 8.2.
Consider the 16 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2 with $w$-eigenvalue 1-1. Let $f_{1},f_{2},f_{3}$ be any three distinct ones among them. Then
${\mathrm{ord}}_{s=2}L\left(s,f_{1}\otimes f_{2}\otimes f_{3}\right)\geq 3.$
1.1 follows from 8.2.